OpenStudy (anonymous):

Simplify, if possible. (b^-2/b^3)^2

OpenStudy (anonymous):

1/(b^25)

OpenStudy (anonymous):

any number raised to a negative exponent can be reciprocated, for example: x^-1 = 1/x and if you have x^-1/x^2, just add the exponents like 1/x^1+2 = 1/x^3 and if you have (1/x^2)^2, just multiply that second exponent through (it needs to be distributed. it'd be 1/x^4.

OpenStudy (anonymous):

\[\left(\begin{matrix}b ^{-2} \\b ^{3}\end{matrix}\right)^{2}=\left( b ^{-2-3} \right)^{2}=\left( b ^{-5} \right)^{2}=b ^{-5\times2}=b ^{-10}\]

OpenStudy (anonymous):

I'm sorry, I squared it by mistake. It should be, 1/(b^10)

OpenStudy (anonymous):

\[\left(\begin{matrix}a ^{m} \\a ^{n}\end{matrix}\right)=a ^{m -n}\]

OpenStudy (anonymous):

\[\left( a ^{m} \right)^{n}=a ^{m.n}\]